Five jars contain a total of 200 pieces of candy. How many pieces of candy are in each jar if each jar contains 2 more pieces than the previous jar?

Found 2 solutions by CubeyThePenguin, greenestamps:
Answer by CubeyThePenguin(3113)

pieces of candy: n, n+2, n+4, n+6, n+8

n + (n+2) + (n+4) + (n+6) + (n+8) = 200
5n + 20 = 200
5n = 180
n = 36
pieces in each jar: 36, 38, 40, 42, 44

Answer by greenestamps(13200)



The numbers form an arithmetic progression, so the third (middle) jar contains the average number of candies of all the jars.

So the "middle" jar has 200/5=40 pieces.

Then since each jar has 2 more than the previous one, the numbers of candies in the jars are 36, 38, 40, 42, and 44.